Cal and subcomponents status overview
This presentation is the property of its rightful owner.
Sponsored Links
1 / 19


  • Uploaded on
  • Presentation posted in: General

CAL and SUBCOMPONENTS STATUS OVERVIEW. T. Sch örner-Sadenius Hamburg University, on behalf of the CAL group ZEUS Meeting DESY 28 February – 4 March 2005. Outline. People and manpower situation CAL status: current status and recent problems CAL DQM development since Padova meeting

Download Presentation


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.

- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript

Cal and subcomponents status overview


T. Schörner-Sadenius

Hamburg University, on behalf of the CAL group

ZEUS Meeting DESY28 February – 4 March 2005



  • People and manpower situation

  • CAL status: current status and recent problems

  • CAL DQM development since Padova meeting

  • Current status of CAL, CFLT(P) and subcomponents(SRTD, SRTDFLT, PRES, F/R PRES, FNC, LED, LASER)

TSS: CAL and Subcomponents

Calorimeter manpower i general coordination

Calorimeter Manpower IGeneral, Coordination

Merging of CAL and HES groups slowly progressing

  • CAL DQM shifters now also do HES DQM checks

  • Progress also on organisational matters, but not on expert side.

    Current CAL coordination

  • TSS: general, DQM, students

  • Wladimir Hain (DESY): general, infrastructure

  • Kerstin Borras (DESY): so far HES general and DQM

    Mara Soares joined the CAL team and will take over CAL DQM coordination from TSS

  • TSS  run coordinator from April on.

  • Mara already now is responsible for a number of DQM tasks.

  • Thanks for joining and good luck!

    New permanent DESY physicist to join CAL team within few months.

Request for oneHES post-doc ordedicated studentto backup Kerstin!


TSS: CAL and Subcomponents

Calorimeter manpower ii

Calorimeter Manpower II

DQM CoordinatorThomas Schörner HH

InfrastructureWladimir Hain DESY

NEVIS and FE Electronics: Bill Schmidke Columbia

Hardware Joker: Cathy Farrow Argonne/DESY

LASER: Roberval Walsh McGill

SRTD: Roberval WalshMcGill

LED: Didar Dobur Freiburg

B-PRES: S. Chekanov + H. Lim Argonne/DESY

SRTDFLT: John Loizides (CFLT) Argonne

F/RPRES, FNC, BCAL basesno dedicated manpower!

Michele SumstineWisconsin (-03/2005)Andree Robichaud-V. McGill (-04/2005)

Ying Cui Toronto

Marcos Jimenez Madrid

Avraam KeramidasNIKHEF

Gabriel GrigorescuNIKHEF

Cathy FarrowArgonne/DESY

Changh Yi ZhouMcGill

General REsponsibilities

Thanks to Michele andAndree and welcome toChangh Yi!

DQM Experts

We would appreciate oneor two new CAL students

TSS: CAL and Subcomponents

Current cal status as of yesterday noon v3959

Current CAL Statusas of yesterday noon: v3959

Two holes: B22, tower 4: UNO problems

B12, tower 8: UNO and QINJ

It is reassuring to know that we are stable around 270 bad channels,if freshly calibrated …

TSS: CAL and Subcomponents

Cal problems during the past few months sometimes overlaid

CAL Problems …… during the past few months – sometimes overlaid

  • Readout failure in december  only cured with help from Antonio P. and Henk B. from Amsterdam after three days of investigation.

  • Disk failure on backup readout computer AJAX and major problems with DQM PCs (not used for online)

  • Timing problems at the CAL-CFLT(P) boundary.

  • BCAL module 16: often many (>30) bad channels, cabling?

  • HV of BCAL module 22 fixed (Wladimir changed HV crate), then module 21 was giving some problems.

  • Os9 hc3control after A. Kotanski’s intervention more or less stable.

Luckily there is always somebody who knows more than I do!Thanks especially to Alessandro and the Nikhef colleagues!

TSS: CAL and Subcomponents

Cal status ctd computing spares

CAL Status ctd. …… Computing, spares …

Spare situation:

  • Front-end cards: C. Farrow efficient in repairing old ones – adequate suppluy of spares

  • NEVIS electronics: spares for all modules; format control card exchanged recently (debugging readout problem).

  • Digital cards: a few spares available; have to investigate 40-50 old, taken-out cards and repair if necessary. Also many ADCs around.

  • FCAL EMC PMTs (bases): 19 spares, could become one critical item.

  • BCAL bases: status unclear – my fault!

  • 9 newly repaired digital crate power supplies just came back!

    Computing situation

  • Teucer and Ajax (Solaris stations) running stably; however, old operating system not supported anymore (v2.4?), and recently disk failure on Ajax. But spares around.

  • 2+1 offline/DQM PCs which also gave problems recently but are stable now.

  • Slightly improving backup situation.

TSS: CAL and Subcomponents

Cal and subcomponents status overview

Bad Channel Historysince Padova week

BC updates

Number of bad channelsis more or less stable with270 or so (after calibration).

1-2 stable holes in BCAL 22and 22.

Some fluctuations inbetween due to localizedproblems, mostly solvedduring access days.


Increase since shutdown

Now stable

Many problems come andgo every now and then;especially in BCAL uncer-tainty of ~100 channels(B16, B29 etc.)

Localizedproblem solvedduring access day.

TSS: CAL and Subcomponents

Hv status since last zeus week influence of beam injection dump on hv settings

R13 t12 EMC1-

R13 t12 EMC1+

F13 t12 EMC1-

F13 t12 EMC1+

R13 t12 EMC2-

R13 t12 EMC2+

F3 t11 EMC1-

F3 t11 EMC1-

HV Status since last ZEUS weekInfluence of beam injection / dump on HV settings

Small trends of the level of 0.1-1% over half a year visible.

TSS: CAL and Subcomponents

Influence of hera on cal beam trigger hermes affect kinematic distributions

Influence of HERA on CALBeam, trigger, HERMES affect kinematic distributions

Sometimes difficult for CAL DQM to judge on data quality.

TSS: CAL and Subcomponents

Cal and subcomponents status overview

CFLT Status

M.Jaworski, D.Reeder, W.Smith

D.Kcira, A.Savin, L.Li

A.Everett, P.Ryan, M.Rosin,

E.Brownson, T.Danielson

CFLT efficiency

RCAL4 Iso_e: 95-97% instead of 98% not a hardware issue – modify DQM toidentify problem

Hardware status

Bad channels

Detector side:2x2 channels ½ height (cable) no 0’s


Two new Alphas are completely installed

Old online machine ZAXP03 still can be used

2 bad serial cables in BCAL:sometimes no shut for ½ module


Offline and online DQM jobs are running

TEC cards: 16 good, 8 bad

Include FCLR and “location check” in the DQM

TSC spare is o.k.

TSS: CAL and Subcomponents

Cal and subcomponents status overview

TSS: CAL and Subcomponents

Cal and subcomponents status overview

Status of FHES


red/blue:’97 installed modules  possibly tiny water leaks

white:’98 installed modules  no indications for water leaks

green:’01/’03 installed modules  no water leaks

FHES 11:

repaired and re-installed in Dec ’04,

end of Jan’05 one half ski (1/6 module) failed electrically  should be repaired in summer break.

  • FHES 9, 10, 13, 15, 16 :

  • monitored closely,

  • all modules recovered, some half skis have high dark current, but are still biased

  • All modules completely working and read out !

    Plan: take out, repair and re-install FHES modules 9, 10, 11, (not 13!), 15, 16 in the summer break.

TSS: CAL and Subcomponents

Cal and subcomponents status overview

Ski – Cooling System:

Status of RHES:

  • no major problem, few trips of HC3 power supplies

  • few failures in half modules due to connection on power distribution card

    • check all modules in the summer break.

Status of DAQ:

  • no major problem

  • run control modified, such that an automatic repetition is initiated in case the SETUP runs into a timeout:

    • much better performance, only two drawbacks to be improved.

  • with the present pump-stand it was possible to lower the under-pressure even further:+0.2bar  -0.2bar

    • under-pressure over almost the full module height

  • several problems: air-leaks to be found, air relief device not working, …

  • anti-corrosive solvent added:

    • lots of problems with foam production blocking flow

      After a quite busy time now finally running stable !


  • Summary

  • lots of troubles with the new cooling settings had to be solved

  • all “possibly leaky” modules in FHES recovered and working 

  • only few failures in the electronics in the meantime 

    HES took good data continuously during the past months 

TSS: CAL and Subcomponents

Srtd status roberval

SRTD Status (Roberval)

What happened since Padova?

Bad channel history since Padova

Current situation:

TSS: CAL and Subcomponents

Cal and subcomponents status overview

S. Chekanov, W. Hain, H. Lim, S. Magill

  • Replaced: Fuses, power crates, readout cables, control cards

  • Problem reproducable in test stand with high temperature: serial data get quickly messed up if fans off!

    • Chip that distributes signals on control car

  • Fan problem? Fans with control signals? Aging of control card chips – replace?

  • DQM was moved to AFS

  • Number of bad HV channels as in 8/2004: 14 (of 416)

TSS: CAL and Subcomponents

Subcomponents status i basically no news

Subcomponents Status I… basically no news!

F/RCAL Presampler (no dedicated manpower):

  • FPRES south: problems in setting HV, no spares

  • FPRES (Wladimir): order of channels wrong in Adamo table? Has anybody ever used this?

  • RPRES okay

  • However, system has no dedicated man power; needs calibration etc.

  • Who is using which part of PRES? Historically only RPRES used?

  • Currently one bad channel.

    FNC (Info from Bill, Wladimir):

  • Test stand working, dubious front-end card needs inspection

  • Some FECs and the delay box need to be checked.

  • Sometimes giving problems in the readout (digital crate) exclude from readout?

    LED for F/RCAL (Freiburg):

  • Not in use currently. Situation rather unclear.

TSS: CAL and Subcomponents

Subcomponents status ii basically no news

Subcomponents Status II… basically no news!

Laser (Roberval Walsh):

  • Not really trusted for calibration/monitoring purposes; however useful for readout debugging in some situations.

  • Currently switched off; some spares should have been replaced ages go. Money? Send to company?

  • Recently problems with LASER test triggers  excluded from test triggers.

    SRTDFLT (John Loizides):

    Situation unchanged; stable.

TSS: CAL and Subcomponents



  • CAL basically running stably with acceptable number of bad channels

    • Status of online and SlowControl systems is satisfactory

    • Calibration procedures are still getting more stable and automatized

  • Some severe problems in the past few months show vulnerability of the system. Sometimes external help necessary!

    • Computing (now maybe better under control)

    • Readout system

  • Manpower situation still satisfactory.

    • We would like 1-2 new CAL students.

    • Looking forward to welcome new DESY scientist in the CAL team.

    • One post-doc / designated student for HES would be appreciated.

    • Some subcomponents uncovered. How much effort can we / do we want to invest for PRES? FNC? LED? LASER?

TSS: CAL and Subcomponents

  • Login